# Understanding Performance of Long-Document Ranking Models through Comprehensive Evaluation and Leaderboarding

### v3

@@ Abstract
We evaluated 20+ Transformer models for ranking of long documents (including recent LongP models trained with FlashAttention) and compared them with a simple FirstP baseline, which applies the same model to the truncated input (at most 512 tokens). We used MS MARCO Documents v1 as a primary training set and evaluated both zero-shot transferred and fine-tuned models.
On MS MARCO, TREC DLs, and Robust04 no long-document model outperformed FirstP by more than 5% in NDCG and MRR (when averaged over all test sets). We conjectured this was not due to models’ inability to process long context, but due to a positional bias of relevant passages, whose distribution was skewed towards the beginning of documents. We found direct evidence of this bias in some test sets, which motivated us to create MS MARCO FarRelevant (based on MS MARCO Passages) where the relevant passages were not present among the first 512 tokens.
Unlike standard collections where we saw both little benefit from incorporating longer contexts and limited variability in model performance (within a few %), experiments on MS MARCO FarRelevant uncovered dramatic differences among models. The FirstP models performed roughly at the random-baseline level in both zero-shot and fine-tuning scenarios. Simple aggregation models including MaxP and PARADE Attention had good zero-shot accuracy, but benefited little from fine-tuning. Most other models had poor zero-shot performance (sometimes at a random baseline level), but outstripped MaxP by as much as 13-28% after finetuning. Thus, the positional bias not only diminishes benefits of processing longer document contexts, but also leads to model overfitting to positional bias and performing poorly in a zero-shot setting when the distribution of relevant passages changes substantially. We make our software and data available. [^1] https://anonymous.4open.science/r/long_doc_rank_model_analysis_v2-78E9/.
Understanding Performance of Long-Document Ranking Models through Comprehensive Evaluation and Leaderboarding
Leonid Boytsovthanks: Work done outside the scope of Amazon employment, primarily while employed by Bosch.thanks: Equal contribution. Amazon AWS AI Labs Pittsburgh, USA leo@boytsov.info                        David Akinpelu 22footnotemark: 2 Louisiana State University Baton Rouge, USA akinpeluakorede01@gmail.com
Tianyi Linthanks: Work done while studying at Carnegie Mellon University.and Fangwei Gao33footnotemark: 3and Yutian Zhao33footnotemark: 3and Jeffrey Huang33footnotemark: 3
Nipun Katyal  and Eric Nyberg Carnegie Mellon University Pittsburg, PA, USA
Figure 1: Average relative gain (in %) vs. relative increase in run-time compared to respective FirstP baselines on MS MARCO, TREC DL 2019-2021, and Robust04 (for a representative subset of models).
Figure 2: Zero-shot vs. fine-tuned performance on MS MARCO FarRelevant for a representative set of models.
@@ 1 Introduction
Transformer models Vaswani et al. (2017)—such as BERT Devlin et al. (2019)—pretrained in a self-supervised manner considerably advanced state-of-the-art of core natural language processing (NLP) Devlin et al. (2019); Radford et al. (2018) and information retrieval Nogueira and Cho (2019). However, due to quadratic cost of the self-attention with respect to an input sequence length, a number of “chunk-and-aggregate” approaches were proposed and evaluated Dai and Callan (2019); MacAvaney et al. (2019); Boytsov and Kolter (2021); Li et al. (2024), but existing studies typically have at least one of the following shortcomings:
  • Reliance only on small-scale query collections such as TREC DL Craswell et al. (2020, 2022), Robust04 Voorhees (2004), and Gov2 Terabyte Clark et al. (2005);
  • Lacking systematic comparison with respective FirstP baselines, which consists in applying the same model to input truncated to the first 512 tokens,
  • Lacking comparison with LongP models—directly supporting long inputs—such as sparse-attention models Longformer and BigBird Beltagy et al. (2020); Zaheer et al. (2020), or more recent full-attention models trained with FlashAttention Dao et al. (2022);
  • Using undisclosed seed-selection strategies, which can restrict reproducibility since there can be substantial (in the order of few %) differences due to using different seeds.
To fill this gap we evaluated over 20 recent models for ranking of long documents and carried out their systematic comparison using two popular document collections: MS MARCO Documents v1/v2 Craswell et al. (2020) and Robust04 Clarke et al. (2004), diverse query sets (both large and small) and multiple training seeds. We found that ranking models capable of processing long documents—including LongP models with sparse or full attention—showed little to no improvement compared to their respective FirstP baselines (which truncated documents to satisfy the input-sequence constraint of most off-the-shelf Transformer models, i.e., 512 tokens).
This finding is generally in line with previously reported results (see § B.4) and an ablation experiment showed that limited improvement over FirstP was not related to the choice of the backbone Transformer model (see Table 7). Furthermore, we used our best models to produce several high-ranking runs on a competitive leaderboard. This, in our view, strengthens the credibility of our evaluation.
From the efficiency-effectiveness plot in Fig. 1, we can see that all long-document models are at least 2$\times$ slower than respective FirstP baselines. The biggest average gain of merely 5% is achieved by the PARADE Attn model (with a BERT-base backbone) at the cost of being 2.5$\times$ slower than its FirstP baseline. All LongP models are even slower and show less improvement. Given such small benefits at the cost of a substantial slow-down, one could question practicality of such models and suggest using FirstP variants instead.
@@
  • RQ1: What is the reason for the lackluster performance of long-document models?
  • RQ2: How much progress has the community made in improving long-document ranking models?
To answer these questions, we started with analyzing a distribution of relevant passages in the MS MARCO document collection and found evidence of a substantial positional bias, namely, relevant passages tended to appear in the beginning of documents. This finding—which partially answers RQ1—prompted an additional research question:
@@
  • RQ3: How robust are long-document models to the positional-bias of relevant passages?
To further support the relevance-bias hypothesis and answer RQ3, we constructed a new synthetic collection MS MARCO FarRelevant where relevant passages were not present among the first 512 tokens. Using MS MARCO FarRelevant, we evaluated zero-shot transferred as well as fine-tuned models and found the following (see Fig. 2):
  • The FirstP models performed roughly at the random-baseline level in both zero-shot and fine-tuning modes (RQ3);
  • Simple aggregation models including MaxP and PARADE Attention had good zero-shot accuracy, but benefited little from fine-tuning on MS MARCO FarRelevant (RQ3);
  • In contrast, other long-document models had poor zero-shot performance (sometimes at a random baseline level), but outstripped respective MaxP baselines by as much as 13.3%-27.7% after finetuning (RQ3);
  • Not only positional bias diminished benefits of processing longer document contexts, but it also lead to models’ overfitting to the bias and performing poorly in a zero-shot setting when the distribution of relevant passages changed substantially (RQ3);
  • Although PARADE Transformer models were more effective than other models on standard collections, their advantage was small (a few %). In contrast, on MS MARCO FarRelevant, PARADE Transformer (ELECTRA) outperformed the next competitor Longformer by 8% and PARADE Max (ELECTRA)—an early chunk-and-aggregate approach—by as much as 23.8% (RQ2).
Our key contributions are as follows:
  • We carried a comprehensive evaluation of 20+ long-document ranking models, which included both the chunk-and-aggregate models as well as the models that directly supported long inputs (using both the standard collections MS MARCO Documents v1/v2 and Robust04 as well as the new synthetic collection MS MARCO FarRelevant);
  • We contributed to the nascent field of analytical experimentation with a full control of outcomes by creating a new dataset MS MARCO FarRelevant, which we made available together with code. [^2] https://anonymous.4open.science/r/long_doc_rank_model_analysis_v2-78E9/.
  • Our study confirmed superiority of PARADE Li et al. (2024) models, but also showed their limited benefits on standard collections, which we attributed to the existence of positional bias of relevant passages (in such collections);
  • We used MS MARCO FarRelevant to support the positional-bias hypothesis as well as to demonstrate that best long-document ranking models substantially (by up to 27.7%) outperform simpler baselines (such as MaxP) when training/fine-tuning data is available. However, they can also suffer more from the distribution shift and perform much worse in the zero-shot scenario.
2 Methods
2.1 Related Work
Neural Ranking models have been a popular topic in recent years Guo et al. (2019), but the success of early approaches was controversial Lin (2019). This changed with an introduction of a bi-directional encoder-only Transformer model BERT Devlin et al. (2019), which was a successor of GPT Radford et al. (2018) and ELMO Peters et al. (2018). BERT was hugely successful and its resounding success can be attributed to a combination of the large model size and massive pretraining using self-supervision. A number of different Transformer models such as ELECTRA Clark et al. (2020), and DEBERTA He et al. (2021) improve upon BERT using different training strategies and/or datasets. However, due to their architectural similarities we—following Lin et al Lin et al. (2021)—collectively call them as BERT models.
Nogueira and Cho were first to apply BERT to ranking of text documents Nogueira and Cho (2019). In the big-data regime—most notably in the TREC deep learning track Craswell et al. (2020)—BERT models outperformed prior neural and non-neural approaches by a large margin. They were also quite successful for several small-scale query collections outperforming previous neural and traditional approaches Li et al. (2024); MacAvaney et al. (2019); Dai and Callan (2019).
Despite their impressive performance, neural models are susceptible to the distribution shift and learning superficial features. Several authors found that neural rankers applied to out-of-domain data do not always outperform BM25 Thakur et al. (2021); Mokrii et al. (2021). They can also be confused by superficial text modifications such as adding distractor sentences MacAvaney et al. (2022). Likewise, ranking performance can decrease if a query is reformulated Penha et al. (2022). Weller et al. Weller et al. (2023) showed that neural models are not effective to “spot” negation and often perform at random level in this respect. However, we are not aware of the prior work systematically studying robustness to positional biases of relevant passages.
The Transformer model Vaswani et al. (2017) uses an attention mechanism Bahdanau et al. (2015) where each sequence position can attend to all the positions in the previous layer. Because self-attention complexity is quadratic with respect to a sequence length, direct processing of long documents is not always practical. Thus, a vast majority of existing Transformer models limit the input length to be at most 512 (subword) tokens.
Until recently, there have been two general approaches to handling long documents: localization of attention and splitting documents into chunks each of which is processed separately. Attention-localization approaches combine a limited-span (i.e., a sliding window) attention with some form of a selective global attention. There are many such approaches proposed (see, e.g., a survey by Tay et al. 2020) and it would be infeasible to evaluate them all. Instead we consider two popular models: Longformer Beltagy et al. (2020) and Big-Bird Zaheer et al. (2020).
With a document-splitting approach, one has to split documents into several chunks, process each chunk separately, and aggregate results, e.g., by computing a maximum or a weighted prediction score Yilmaz et al. (2019); Dai and Callan (2019). With respect to training approaches, the MaxP and SumP models by Dai and Callan Dai and Callan (2019) assume that each chunk in a relevant document is relevant. However, this assumption is problematic as the degree of relevance varies from passage to passage. Yilmaz et al. Yilmaz et al. (2019) work around this problem by training a MaxP BERT model on short documents and zero-transfer it to long documents. In this study we work around this problem by training all document-splitting approaches including MaxP Dai and Callan (2019) in the end-to-end fashion, i.e., by plugging aggregated document-level scores directly into a loss function (analogous to training of CEDR MacAvaney et al. (2019) and PARADE Li et al. (2024) models).
More recently, it has also become possible to train longer-context models using FlashAttention Dao et al. (2022). FlashAttention computes attention exactly and it does not eliminate quadratic complexity. However, it dramatically speeds ups training while reducing memory requirements by using an IO-efficient computation approach.
Because our primary focus is accuracy and we aim to understand the limits of long-document models, we exclude from evaluation several recent models (e.g., Hofstätter et al. (2021); Zou et al. (2021)) that achieve better efficiency-effectiveness trade-offs by pre-selecting certain document parts and feeding only selected parts into a BERT ranker.
Recently, several teams have focused on creating challenging benchmarks for long-document retrieval. A recent LoCo v1 Saad-Falcon et al. (2024) benchmark has 12 datasets. Despite 11 out of 12 collections has average document lengths in the order of dozens of thousands tokens, the E5 model with a 512 token input limit achieves high NDCG@10 scores (in the range of 0.4-0.85) for seven out of 12 LoCo v1 datasets. This prompted Zhu et al., 2024 to propose a more challenging LongEmbed benchmark containing a mix of real and synthetic datasets Zhu et al. (2024).
Table 1: Distribution of Start/End Positions of Relevant Passages Inside Documents
input chunk #
MS MARCO dev
FIRA
Hofstätter et al. (2020b)
(crowd-sourced)
start end start end
1 85.9% 71.0% 83.8% 76.4%
2 9.1% 14.9% 9.9% 15.3%
3 2.6% 6.1% 2.3% 3.9%
4 1.2% 3.0% 2.2% 2.2%
5 0.6% 1.4% 0.7% 0.9%
6 0.6% 1.2% 0.4% 0.5%
6+ 0.1% 2.5% 0.7% 0.7%
  • Chunk size is 477 BERT tokens.
Table 2: Document Statistics
data set # of documents average # of
BERT tokens
per document
MS MARCO v1 3.2M 1.4K
MS MARCO v2 12M 2K
Robust04 0.5M 0.6K
MS MARCO FarRelevant 0.53M 1.1K
Table 3: Query Statistics
# of queries
avg. # of
BERT tokens
avg. # of
pos. judgements
MS MARCO v1
MS MARCO train 352K 7 1
MS MARCO dev 5193 7 1
TREC DL 2019 43 7 153.4
TREC DL 2020 45 7.4 39.3
MS MARCO v2
TREC DL 2021 57 9.8 143.9
Robust04
title 250 3.6 69.6
description 250 18.7 69.6
MS MARCO FarRelevant
train 50K 7.0 1
test 1K 7.0 1
2.2 Data
Our primary datasets include two MS MARCO Documents collections (v1 and v2) Bajaj et al. (2016); Craswell et al. (2020, 2022), Robust04 Voorhees (2004), and associated query sets. In addition, we created a collection MS MARCO FarRelevant by using passages and relevance judgments from the MS MARCO Passages collection.
Robust04 is a small collection of 0.5M documents that has a mixture of news articles and government documents some of which are quite long. Yet it has only a small number of queries (250), which makes it a challenging benchmark for training models in a low-data regime. Each query has a title and a description, which represent a brief information need and a more elaborate request (often a proper English prose), respectively. We use Robust04 in a cross-validation settings with folds established by Huston and Croft Huston and Croft (2014) provided via IR-datasets MacAvaney et al. (2021). All datasets are in English. Document and query statistics are summarized in Tables 2 and 3.
MS MARCO v1 was created from the MS MARCO reading comprehension dataset Bajaj et al. (2016) and it has two related collections: passages and documents. MS MARCO v1 comes with large query sets, which is particularly useful for training and testing models in the big-data regime. These query sets consist of question-like queries sampled from the Bing search engine log with subsequent filtering Craswell et al. (2020). Note that queries are not necessarily proper English questions, e.g., “lyme disease symptoms mood”, but they are answerable by a short passage retrieved from a set of about 3.6M Web documents Bajaj et al. (2016).
MS MARCO v1 test sets were created in two stages, where initially relevance judgements were created for the passage variant of the dataset. Then, document-level relevance labels were created by transferring passage-level relevance to original documents from which passages were extracted. To assess positional bias, we mapped relevant passages (from the MS MARCO Passage collection) to their positions in documents. Because document and passage texts were collected at different times this lead to some content divergence Craswell et al. (2020) and made exact mapping impossible: In particular, Hofstätter et al. 2020b were able to match only 32% of the passages:
We deemed such mapping insufficient: To obtain a more comprehensive mapping we resorted to approximate matching and were able to match about 85% of the passages. We manually inspected a sample of matched passages to ensure that the matching procedure was reliable. Moreover, the distribution of positions of relevant passages matched that of a related FIRA dataset Hofstätter et al. (2020b), where such information was collected by crowdsourcing. Positional bias information is summarized in Table 1.
Relevance labels in the training and development sets are “sparse”: There is about one positive example per query without explicit negatives. In addition to sparse relevance judgements—separated into training and developments subsets—there is a small number (98) of queries that have “dense” judgements provided by NIST assessors for TREC 2019 and 2020 deep learning (DL) tracks Craswell et al. (2020).
MS MARCO v2 collections was created for TREC 2021 DL track. It is an expanded version of MS MARCO v1 and uses a subset of sparse relevance judgements from MS MARCO v1. In the training set, newly added documents do not have any (positive or negative) judgments, which created a bias and made MS MARCO v2 training set less useful than that of MS MARCO v1.
The MS MARCO FarRelevant collection was created from the MS MARCO passage collection in such a way that each document contained exactly one relevant passage and this passage did not start before token 512 (see algorithm in the Appendix § B.1). Moreover, we created just a single relevant document for each training or testing query. MS MARCO FarRelevant is a variant of a the needle-in-the-haystack test Saad-Falcon et al. (2024); Zhu et al. (2024). It is designed to be textually similar to MS MARCO Documents but with different positional biases for relevant passages. Due MS MARCO having a non-commercial license, MS MARCO FarRelevant has the same licensing restriction.
Although we generated about 7K test queries and about 500K training queries, we used only 50K and 1K queries for fine-tuning and testing, respectively. On one hand, this was sufficient for accurate training and testing and, on the other hand, it reduced experimentation time and cost.
2.3 Overview of Selected Methods
Due to space constraints, a detailed description is given in the Appendix § A. In summary, all methods can be divided into split-and-aggregate (SplitP) methods and LongP methods that “natively” support longer documents inputs. SplitP use either simple aggregating operations (averaging, summing, taking the maximum) or an aggregator neural network. CEDR MacAvaney et al. (2019), PARADE Attention Li et al. (2024), and Neural Model 1 Boytsov and Kolter (2021) aggregate using simple neural networks, whereas PARADE Transformer models aggregator is a smaller Transformer Li et al. (2024).
We focused on cross-encoding rankers, which process queries concatenated with documents Nogueira and Cho (2019). As a reference point we also tested a bi-encoding E5-4K model, which had strong performance on LongEmbed benchmark with context sizes under 4K tokens Zhu et al. (2024). E5-4K was tested as a ranking model and only in the zero-shot mode (without fine-tuning).
Nearly all rankers use only BERT models (i.e., bi-directional encoder-only Transformers) and have in total 100M-200M parameters (see Table 6). In addition, inspired by a recent success of LLM-rankers Pradeep et al. (2023); Ma et al. (2023), we tested a much larger cross-encoding decoder-only (“causal”) Transformer model. Specifically we chose a 1B-parameter TinyLLAMA model due to its impressive performance for its relatively small size Zhang et al. (2024).
Table 4: Ranking Performance on MS MARCO, TREC DL, and Robust04.

### v4

# Positional Bias in Long-Document Ranking: Impact, Assessment, and Mitigation

@@ Abstract
We tested over 20 Transformer models for ranking long documents (including recent LongP models trained with FlashAttention and RankGPT models “powered” by OpenAI and Anthropic cloud APIs). We compared them with the simple FirstP baseline, which applied the same model to truncated input (up to 512 tokens). On MS MARCO, TREC DL, and Robust04 no long-document model outperformed FirstP by more than 5% (on average). We hypothesized that this lack of improvement is not due to inherent model limitations, but due to benchmark positional bias (most relevant passages tend to occur early in documents), which is known to exist in MS MARCO. To confirm this, we analyzed positional relevance distributions across four long-document corpora (with six query sets) and observed the same early-position bias. Surprisingly, we also found bias in six BEIR collections, which are typically categorized as short-document datasets. We then introduced a new diagnostic dataset, MS MARCO FarRelevant, where relevant spans were deliberately placed beyond the first 512 tokens. On this dataset, many long-context models—including RankGPT—performed at random-baseline level, suggesting overfitting to positional bias. We also experimented with debiasing training data, but with limited success. Our findings (1) highlight the need for careful benchmark design in evaluating long-context models for document ranking, (2) identify model types that are more robust to positional bias, and (3) motivate further work on approaches to debias training data. We release our code and data to support further research.
(a) Robust04
(b) ClueWeb12 (WebTrack)
(c) TREC DL 2019-2021 (combined)
Figure 1: Positional relevance bias for three long-document collections (best viewed in color). We show a distribution of first relevant passage positions (red bars) vs. relevant document lengths (blue bars). Lengths and offsets are measured in the number of subword tokens (BERT-base tokenizer). See more results (including BEIR) in Figures 7 and 8 in Appendix C.1.
@@ 1 Introduction
Various advances in Transformer architectures—including sparse attention Zaheer et al. (2020); Beltagy et al. (2020) and FlashAttention Dao et al. (2022)—have motivated a growing interest in long-document ranking and retrieval. However, despite the ability of these models to process substantially more text, on popular retrieval benchmarks the improvements by these models over simpler truncation-based approaches remain surprisingly modest Dai and Callan (2019); Gao and Callan (2022); Boytsov et al. (2022); Coelho et al. (2024). A widely used truncation-based FirstP baseline Dai and Callan (2019)—where models score only the first 512 tokens of each document—often performs competitively or sometimes even better than long-context counterparts (see, e.g., Table 1).
Despite anecdotal knowledge about the presence of this phenomenon in the MS MARCO document-retrieval collection among TREC Deep Learning track participants and in some early reports Hofstätter et al. (2020b, 2021a); Boytsov et al. (2022) the available evidence has been scattered and incomplete. In particular, despite the track’s five-year history, none of the track’s overview papers have mentioned this issue Craswell et al. (2020, 2021a, 2021b, 2022, 2023).
Moreover, it remains unclear whether these limitations stem from model deficiencies or from characteristics of the benchmarks themselves. In this paper, we initially hypothesized that both factors—model robustness and benchmark design—may be responsible for the limited gains achieved by long-document models over FirstP baselines. However, our findings suggest that benchmark design, particularly positional relevance bias, is the dominant factor.
To verify our research hypotheses, we first conducted a large-scale, systematic study of over 20 Transformer-based ranking models Devlin et al. (2019); Vaswani et al. (2017) for long-document retrieval. This was done using three popular document collections—MS MARCO Documents v1/v2 Craswell et al. (2021a) and Robust04 Clarke et al. (2004)—along with diverse query sets (both large and small), several Transformer backbones, and multiple training seeds. In addition to locally trained models, we also assessed a listwise LLM ranker RankGPT Sun et al. (2023) “powered” by OpenAI OpenAI (2023) and Anthropic Anthropic (2024) cloud APIs.
Despite the increased context capacity of long-document models, we found that none of them consistently outperformed their FirstP baselines by more than 5% on average. An ablation experiment showed that limited gains over FirstP were not attributable to the choice of a Transformer backbone model (see Table 5).
Next, we estimated positional relevance bias across five document collections and more than six query sets. As can be seen from Fig. 1, in the vast majority of cases, the first relevant passage occurred within the initial 512 tokens. In contrast, the distribution of relevant passage positions is more uniform, with a pronounced long tail. This confirms that positional bias is substantial not only in MS MARCO, but also in other TREC collections. Surprisingly, we also found bias in several BEIR collections Thakur et al. (2021), which are typically categorized as short-document datasets (a complete set of plots is provided in Fig. 7 and 8 of Appendix § C.4).
@@
  • RQ1: How robust are long-document models to the positional bias of relevant passages?
[…]
To answer these questions, we constructed a new diagnostic synthetic collection MS MARCO FarRelevant where relevant passages were not present among the first 512 tokens. On this dataset, many long-context models—including RankGPT Sun et al. (2023)—failed to generalize and performed at a random baseline level, suggesting overfitting to positional bias. Poor performance of models on MS MARCO FarRelevant prompted another important question RQ3: Can debiasing of training data mitigate model overfitting to positional bias? We addressed this question by evaluating an existing debiasing approach by Hofstätter et al. (2021a).
Our paper makes the following contributions:
  • We re-examined the issue of positional relevance bias, gathered extensive evidence confirming its presence in both long- and short-document datasets, showed that it negatively affects all models including recent LLM-based rankers, and evaluated the robustness of ranking models against this bias using a new diagnostic dataset MS MARCO FarRelevant.
  • Our work highlights the need for careful benchmark design in evaluating long-context models for document ranking, so that benchmarks do not mask the benefits of long-context models, and identify model types that are more robust to positional bias;
  • We performed an extensive reproduction study of over 20 ranking models using established benchmark collections for long-document retrieval and ranking: MS MARCO Documents v1 and v2 Craswell et al. (2021a) and Robust04 Clarke et al. (2004);
  • We experimented with an existing approach to debiasing training data Hofstätter et al. (2021a) and motivated further research in this area.
Our code and data are available. [^1] https://github.com/searchivarius/long_doc_rank_model_analysis_v2/.
2 Related Work
Neural Ranking models have been a widely studied topic in recent years Guo et al. (2019), though the success of early approaches was debated Lin (2019). This changed with the introduction of BERT, a bidirectional encoder-only Transformer model Devlin et al. (2019), which significantly outperformed previous methods in both NLP Devlin et al. (2019) and information retrieval (IR) tasks Nogueira et al. (2019); Craswell et al. (2021a).
Several Transformer-based models, such as ELECTRA Clark et al. (2020) and DEBERTA He et al. (2021), have improved upon BERT through different training strategies and datasets. However, due to their architectural similarities, we—following Lin et al. (2021)—refer to these collectively as BERT models.
Despite their strong performance, neural models are vulnerable to distribution shifts, often relying on superficial features and exhibiting various biases. They do not consistently outperform BM25 on out-of-domain data Mokrii et al. (2021); Thakur et al. (2021), can be misled by minor text modifications and distractor sentences MacAvaney et al. (2022), or reformulated queries Penha et al. (2022). They also struggle to effectively utilize information located in the middle of long input contexts, in particular for retrieval-augmented generation Liu et al. (2024). Recent work showed that this can be an intrinsic architectural artifact such that the model gives disproportionately high attention to tokens at the beginning or end of the input context, regardless of their actual relevance, which can be partially mitigated by re-calibration Hsieh et al. (2024). However, An et al. (2024) argue that architectural biases can be exacerbated by insufficient explicit supervision during long-context training, which causes models to assume that important information is concentrated near context-window edges.
A related but distinct issue—positional relevance bias—which this study attributes primarily to the characteristics of supervised training data, has been identified in information-retrieval settings, particularly within the MS MARCO document-retrieval collection Hofstätter et al. (2020b); Boytsov et al. (2022); Coelho et al. (2024). Some studies have reported the strong performance of FirstP baselines on long-document retrieval collections, interpreting this as evidence of benchmark-induced positional bias Boytsov et al. (2022); Zhu et al. (2024); Rau et al. (2024). Note, however, that strong performance of FirstP baseline is only indirect evidence, potentially resulting from implementation bugs, suboptimal training methods, or model-inherent biases.
The earliest study on this topic obtained direct evidence of bias only for a small set of TREC DL queries, without studying how bias affects model performance Hofstätter et al. (2020b). Boytsov et al. (2022) tested a variety of long-document ranking models (on MS MARCO and Robust04 datasets) and found them to be only marginally better than the FirstP baseline, which they attributed to positional relevance bias. Coelho et al. (2024) found that two embedding models trained on MS MARCO “dwell” in the beginning and are less effective when relevant information is present elsewhere in a document. These studies, however, (1) had limited evidence that bias was attributable to data, (2) did not evaluate key modern long-context models, and (3) did not assess a debiasing strategy.
Because positional relevance bias can “obscure” benefits of long-document ranking models, Rau et al. (2024) proposed to compare these models with RandP baselines, which score a randomly selected passage. However, we believe this approach is problematic for two reasons. First, since a RandP model often fails to score an entire relevant passage, it artificially underestimates a model’s accuracy, making comparisons with RandP unfair. Second, this approach does not address the core problem of biased benchmarks, which still allow models to exploit the shortcut of focusing only on the document’s initial portion.
Interestingly, Rau et al. (2024) argue that positional bias is strong only in MS MARCO, but not in Robust04. We believe this is because they only compared FirstP and MaxP models in a zero-shot setting (trained on MS MARCO and tested on Robust04), whereas we fine-tuned all models on Robust04 as well. The relatively strong performance of MaxP compared to FirstP in their study may be due to the generally stronger zero-shot performance of MaxP models. We observe similarly small gains over FirstP for both MS MARCO and Robust04, and our passage-relevance analysis confirms that they have comparable levels of positional bias. Thus, despite their widespread use, Robust04 and MS MARCO are not particularly useful for benchmarking of long-document models.
To address the issue with existing benchmarks, Zhu et al. (2024) proposed a LongEmbed benchmark with two synthetic tasks where relevant mini-passages were scattered uniformly across documents with lengths varying from 256 to 32768 tokens. However, as discussed in §C.3, these synthetic sets are quite unnatural and lack diversity. Furthermore, Zhu et al. (2024) provide only small query sets and no in-domain training data, making it difficult to assess the upper performance bound that models can achieve on this dataset. As another important limitation Zhu et al. (2024), only explored training-free extensions of positional encoding and did not investigate methods to debias training data. In contrast, Hofstätter et al. (2021a) proposed to debias training data using a simple yet effective approach. However, they did not evaluate it on challenging long-document datasets.
In this paper, we aim to provide stronger evidence through collecting more comprehensive statistics of relevant passages in documents, experimenting with a diverse set of models, and applying the diagnostic synthetic set, which also demonstrates models’ abilities to adapt to different positional biases.
Due to the quadratic complexity of the Transformer’s attention mechanism Vaswani et al. (2017); Bahdanau et al. (2015), early Transformer models restricted input length to a maximum of 512 (subword) tokens. Until around 2022, two main strategies were used to process long documents: (1) localizing attention and (2) splitting documents into smaller, independently processed chunks. Attention-localization methods apply a limited-span (sliding window) attention and selective global attention. Given the vast number of such approaches (see Tay et al. 2020), evaluating all of them is impractical. Therefore, we focus on two popular models: Longformer Beltagy et al. (2020) and BigBird Zaheer et al. (2020). More recently, it has also become feasible to train long-context models with an IO-efficient FlashAttention algorithm without sparsifying attention Dao et al. (2022).
In summary, methods to tackle longer documents are divided into LongP methods—where longer document are “natively” supported and SplitP methods—where a longer document cannot be processed as a whole and needs to be processed in chunks. The results of each chunk are aggregated together using various aggregation techniques, including computation of a maximum or a weighted-sum prediction score Yilmaz et al. (2019); Dai and Callan (2019); MacAvaney et al. (2019). This includes MaxP Dai and Callan (2019), AvgP, SumP MacAvaney et al. (2019), as well as PARADE Avg and PARADE Max models MacAvaney et al. (2019). MaxP is an important baseline that computes relevance scores for each chunk independently and takes their maximum.
Some SplitP approaches aggregate using simple neural networks. This includes all CEDR MacAvaney et al. (2019) models, the Neural Model 1 Boytsov and Kolter (2021), and the PARADE Attention model Li et al. (2024). In contrast, PARADE Transformer Li et al. (2024) models’ aggregator network is an additional Transformer model. Due to space constraints, a detailed description of document-splitting (SplitP) approaches is provided in the Appendix § D.
The recent success of decoder-only models—commonly known as LLMs—has led to a new generation of cross-encoding LongP models that natively support longer contexts. First, a pre-trained cross-encoding decoder-only model can be directly fine-tuned on a ranking or embedding task Ma et al. (2023). Second, the RankGPT approach Sun et al. (2023) formulates document ranking as a generation task: The model is prompted with a list of documents and an instruction to generate their ranking—an ability made possible through instruction-tuning and/or alignment Wei et al. (2022); Ouyang et al. (2022). When the combined length of concatenated documents exceeds the input context size, RankGPT employs an overlapping sliding window strategy, followed by aggregation of the results.